OpenStudy (anonymous):

Sam did 480 J of work when moving a bag with a mass of 38 kg. He did the work in 4 minutes. What was the power output? 442 W 20 W 120 W 2 W

OpenStudy (anonymous):

Power is Joules/sec. Translate 4 mins to seconds.

OpenStudy (anonymous):

p=w/t where w=work in J t=time in sec p=480/(4*60) w

OpenStudy (anonymous):

soooo 2w?

OpenStudy (anonymous):

Yep.